OpenStudy (anonymous):

omg i need a lot of help what is 1/4 + w =6 1/3

OpenStudy (alexandervonhumboldt2):

first subtract 1/4 from both sides: \[\frac{ 1 }{ 4 }-\frac{ 1 }{ 4 }+w=6\frac{ 1 }{ 3 }-\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

so 1

OpenStudy (alexandervonhumboldt2):

we get \[w=6\frac{ 1 }{ 3 }-\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

um ok i dont even hava a clue still

OpenStudy (alexandervonhumboldt2):

now equal the denominators: \[w=6\frac{ 4 }{ 12 }-\frac{ 3 }{ 12 }\]

OpenStudy (anonymous):

12?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

\[w=6\frac{ 4-3 }{ 12 }=6\frac{ 1 }{ 12 }\]

OpenStudy (anonymous):

ok whats next